OpenStudy (anonymous):

Suppose f(π/6)=7 and f′(π/6)=−4 , and let g(x)=f(x)cosx and h(x)=sinx/f(x) g'(π/6)=? h'(π/6)=?

OpenStudy (mertsj):

\[g(x)=f(x)\cos (x)\] \[g'(x)=f'(x)\cos (x)-f(x)\sin (x)\] \[g'(30)=f'(30)\cos(30)-f(30)\sin(30)\]

OpenStudy (mertsj):

\[g'(30)=-4(\frac{\sqrt{3}}{2})-7(\frac{1}{2})\]

OpenStudy (mertsj):

\[h(x)=\frac{\sin x}{f(x)}\] \[h'(x)=\frac{f(x)\cos x-\sin xf'(x)}{[f(x)]^{2}}\ ]\[h'(30)=\frac{f(30)\cos (30)-\sin (30)f'(30)}{[f(30]^{2}}\]

OpenStudy (mertsj):

\[h'(30)=\frac{f(30)\cos (30)-\sin (30)f'(30)}{[f(30)]^{2}}\]

OpenStudy (mertsj):

\[h'(30)=\frac{7(\frac{\sqrt{3}}{2}-\frac{1}{2}(-4)}{7^{2}}\]

OpenStudy (mertsj):

\[h'(30)=\frac{14\sqrt{3}+2}{49}\]

OpenStudy (mertsj):

I used 30 instead of pi/6 so I wouldn't have to type all those fractions.